OpenStudy (anonymous):

Given the number, two-fifths, generate its equivalent forms as a fraction, a decimal, and a percent. In addition, give a real world example in which each form might be used. use your one words

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@Nnesha @ksardar828 @Michele_Laino

OpenStudy (michele_laino):

\[\frac{ 2 }{ 5 }=\frac{ 2*2 }{ 5*2 }=\frac{ 4 }{ 10 }=0.4\] furthermore: \[\frac{ 2 }{ 5 }=\frac{ 2*2 }{ 5*2 }=\frac{ 4 }{ 10 }=\frac{ 4*10 }{ 10*10 }=\frac{ 40 }{ 100 }=\] =40%

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

thanks!